Use the given degree of confidence and sample data to find a confidence interval for the population standard deviation σ. Assume that the population has a nomn to the same number of decimal places as the sample standard deviation. al distribution. Round the confidence interval limits 7) The mean replacement time for a random sample of 20 washing machines is 10.9 years and the standard deviation is 2.7 years. Construct a 99% confidence interval for the standard deviation, o, of the replacement times of all washing machines of this type. 7
Answer #1

Degrees of freedom df = n-1 = 20 - 1 = 19

99% confidence interval for \sigma is

Sqrt [ (n-1) S2 / \chi ^{2}\alpha/2 ] < \sigma < Sqrt [ (n-1) S2 / \chi ^{2}1-\alpha/2 ]

Sqrt [ 19 * 2.72 / 38.582 ] < \sigma < sqrt [ 19 * 2.72 / 6.844 ]

1.9 < \sigma < 4.5

99% CI for \sigma is ( 1.9 , 4.5 )